OpenStudy (anonymous):

Simplify the expression: cot x sin x - sin quantity pi divided by two minus x. + cos x

OpenStudy (anonymous):

\[\frac{ \cot(x)\sin(x)-\sin(\pi) }{ 2-(x+\cos(x)) }\] Is this the expression?

OpenStudy (anonymous):

Uhm no the expression is\[\cot \left( x \right)\sin \left( x \right)-\sin (\left( \frac{ \pi }{ 2} \right) -x )+\cos \left( x \right)\]

OpenStudy (anonymous):

@VeritasVosLiberabit